Nerve impulses are carried along axons, the elongated fibers that transmit neural signals. We can model an axon as a tube with an inner diameter of 10 μm. The tube wall is insulated, but the fluid inside it has a resistivity of 0.50 Ω⋅m.

Part A

What is the resistance of a 6.0-mm-long axon?

Express your answer with the appropriate units.
Answer #1

here,

the inner diameter , d = 10 um = 10 * 10^-6 m

radius , r = d/2 = 5 * 10^-6 m

resistivity , p = 0.5 ohm.m

length , l = 6 mm = 0.006 m

the resistance , R = p * l /area

R = p * l /( pi * r^2)

R = 0.5 * 0.006 /(pi * ( 5 * 10^-6)^2) ohm

R = 3.82 * 10^7 ohm (Answer)

the resistance of wire is 3.82 * 10^7 ohm

I hope help you !!